A gas in piston-cylinder assembly undergoes a polytrophic expansion, the initial pressure is 3 bar. initial volume is 0.1m3(cube) and the final volume is 0.2m3. determine the work done in kilojoules if
1. N =1.5
2.N=1
3.N=0
1
Expert's answer
2019-12-10T09:55:40-0500

Solution.

1.n=1.5

"W = \frac{p2 \times V2 - p1 \times V1}{1-n}"

"p1 \times V1^n = p2 \times V2^n"

"p2 = p1 \times (\frac{V1}{V2})^n = 1.06 \ bar"

W = 17.6 kJ

2.n=1.0

pV = const

"W = p1 \times V1 \times ln(\frac{V2}{V1})"

W = 20.79 kJ

3.n=0

If n=0, p = const and,

"W = p \times (V2-V1)"

W = 30 kJ

Answer:

1.W = 17.6 kJ

2.W = 20.79 kJ

3.W = 30 kJ